If the sound intensity 4.0 m from a point sound source is 1.0 × 10⁶ Wm², what will be the sound intensity at each of the following distances from the source:


A. 1.0 m? And


B. 2.0 m?



1
Expert's answer
2021-11-17T10:40:06-0500

"I(r) =\frac{P}{4\pi r^2}"

"I(4) = \frac{P}{64\pi }=1.0*10^6\frac{W}{m^2}"

"I(1)= \frac{P}{4\pi }=\frac{I(4)}{16}=0.625*10^5\frac{W}{m^2}"

"I(2)= \frac{P}{16\pi }=\frac{I(4)}{4}=0.25*10^6\frac{W}{m^2}"

"\text{Answer: }"

"A.\ 0.625*10^5\frac{W}{m^2}"

"B.\ 0.25*10^6\frac{W}{m^2}"